OpenStudy (poofypenguin):

If cos(t)=−9/11 where pi

OpenStudy (anonymous):

Find first sin(t) \[\sin(t) = \sqrt{1 - \cos^2(t)}\] Now find : \[\sin(2t) = 2 \sin(t) \cdot \cos(t)\]

OpenStudy (anonymous):

For cos(2t) use : \[\cos(2t) = 2\cos^2(t) - 1\]

OpenStudy (poofypenguin):

how do you find sin(t) in exact value and no decimals?

OpenStudy (anonymous):

We can find it.. Do not simplify the square root;; Like : \[\sin(t) = \sqrt{1 - \frac{81}{121}} \implies \sin(t) = \sqrt{\frac{40}{121}} \implies \sin(t) = \frac{2 \sqrt{10}}{11}\] As interval is given as pi<t<(3pi/2), here sin(t) is negative so: \[\sin(t) = -\frac{2 \sqrt{10}}{11}\]

OpenStudy (poofypenguin):

I understand it so far! Thanks for explaining! So, how would i go about solving the half ones?

OpenStudy (anonymous):

\[\cos(\frac{t}{2}) = 2\cos(t) - 1\]

OpenStudy (poofypenguin):

I've never seen that identity before...is there one for sine too?

OpenStudy (anonymous):

It is double angle formula only..

OpenStudy (anonymous):

I think I forgot the square there sorry...

OpenStudy (ajprincess):

 Wont it be this @waterineyes?